You are attempting to value a call option with an exercise price of $100 and one year to expiration. The underlying stock pays no dividends, its current price is $100, and you believe it has a 50% chance of increasing to $124 and a 50% chance of decreasing to $76. The risk-free rate of interest is 8%. Calculate the call option's value using the two-state stock price model.
Answer #1
High price=Current price*up move=100*1.2667=126.67
Low price=Current price*down move=100*0.8933=89.33
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.08*1)-0.8933)/(1.2667-0.8933)=0.5
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(124-100,0)
=Max(24,0)
=24
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(76-100,0)
=Max(-24,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.08*1)*(0.5*24+(1-0.5)*0)
=11.08